OpenStudy (anonymous):

Find both Unit rates. 30 white cars for every 12 red cars.

OpenStudy (anonymous):

It's asking you for a rate, that means you divide them, it should look something like this\[\frac{ 30white }{ 12red }\]just simplify that to get how many whites there are for every red car, it works the other way as well\[\frac{ 12red }{ 30white }\]will tell you how any red cars there are for every white car

OpenStudy (anonymous):

how do I simplify I mean I have had it explained to me but I still dont understand it @Embryo

OpenStudy (anonymous):

simplifying means, what do both the top and the bottom have in common? Well,\[\frac{ 6*5 }{ 6*2 }\]you have a 6 on both the top and the bottom right? that means they equal 1, because 6 divided by 6 is 1, so that means\[\frac{ 6*5 }{ 6*2 }=\frac{ 5 }{ 2 }\]and that's the simplified rate for white cars for every red cars, it means that you have 5 white cars for every 2 red cars, just work on the other one now